OpenStudy (anonymous):

write a quadratic equation for a vertex at (-3,5) that passes through (4,-1)

OpenStudy (anonymous):

First start fitting the info together the standard form for quadratic equation is like this \(y=(x-h)^2+k\) where (h,k) is the vertex

OpenStudy (anonymous):

y=a(x-h)^2+k i forgot a so put in that vertex you have and try to find what is a from plug in the point (4.-1)

OpenStudy (anonymous):

your equation would be like \(y=a(x+3)^2+5\)

OpenStudy (anonymous):

now all you need to do is find a

OpenStudy (anonymous):

thanks!